# VINP: Variational Bayesian Inference with Neural Speech Prior for Joint ASR-Effective Speech Dereverberation and Blind RIR Identification

### v1

@@ IV-C2 Blind RIR Identification
Unprocessed | 1.48 | 0.70 | 2.37 | 3.20 | 12.7 | 5.6 | 4.7 | 1.31 | 2.82 | 21.1 | 7.7 | 5.6 |
Oracle | - | - | 3.76 | 3.90 | 7.4 | 4.5 | 4.1 | - | - | - | - | - |
GWPE [4] | 1.57 | 0.72 | 2.41 | 3.22 | 11.9 | 5.4 | 4.7 | 1.43 | 2.83 | 17.7 | 7.0 | 5.5 |
SkipConvNet [57] | 2.12 | 0.81 | 3.20 | 3.60 | 12.9 | 6.4 | 5.3 | 2.74 | 3.32 | 26.1 | 9.3 | 7.4 |
CMGAN [23] | 2.85 | 0.90 | 3.82 | 3.81 | 9.2 | 5.1 | 4.5 | 3.86 | 4.00 | 12.5 | 5.8 | 5.2 |
StoRM [25] | 2.34 | 0.86 | 3.73 | 3.96 | 11.4 | 6.0 | 5.1 | 3.72 | 4.01 | 16.8 | 9.9 | 9.7 |
@@
oSpatialNet* [53] | 2.87 | 0.92 | 3.54 | 3.88 | 8.7 | 4.8 | 4.3 | 3.39 | 3.86 | 10.3 | 5.3 | 4.5 |
VINP-TCN+SA+S (prop.) | 2.51 | 0.87 | 3.47 | 3.88 | 8.8 | 5.1 | 4.4 | 3.17 | 3.77 | 11.5 | 6.2 | 5.1 |
VINP-oSpatialNet (prop.) | 2.82 | 0.90 | 3.48 | 3.86 | 8.1 | 4.8 | 4.3 | 3.32 | 3.80 | 9.1 | 5.0 | 4.3 |
@@ IV-E1 Speech Dereverberation
Setting the number of VEM iterations to 100, the dereverberation results are presented in Table II, where the bold font denotes the best two results.
As the metrics of unprocessed speech and oracle speech indicate, a larger ASR system implies a stronger recognition ability for oracle speech and stronger robustness to reverberation. Compared with using a large ASR system alone, introducing an ASR-effective front-end speech dereverberation system can significantly improve the performance of ASR with fewer parameters and lower computational costs. Among the comparison methods, only a few approaches are proven to be effective for ASR, such as GWPE, CMGAN, oSpatialNet*, and the proposed VINP-TCN+SA+S and VINP-oSpatialNet. The classical GWPE yields a marginal improvement in ASR performance. Among the DL-based methods, SkipConvNet and TCN+SA+S encounter failures in ASR. Meanwhile, StoRM demonstrates effectiveness solely when the ’tiny’ ASR model is employed. The reason for such ASR performance lies in the artifact induced by DNN, which is a common phenomenon in similar DL-based single-channel systems [26, 27]. Different from the comparison methods, VINP employs a linear CTF filtering process to model the reverberation effect and leverages the DNN output as a prior distribution of anechoic source speech. During the VBI procedure, VINP takes into account both this prior information and the observed data. By doing this, VINP circumvents the direct utilization of DNN output and thereby reduces the artifacts. Meanwhile, VINP still utilizes the powerful non-linear modeling ability of DNN. Consequently, VINP exhibits a remarkable superiority over the other approaches in ASR performance. When comparing VINP-TCN+SA+S with its backbone network TCN+SA+S, a gap in WER can be discerned, indicating that VINP can make an ASR-ineffective DNN effective. Both oSpatialNet* and VINP-oSpatialNet achieve ASR performance close to that of oracle speech on REVERB SimData when using the small and medium ASR systems. In other situations, a gap still exists, indicating that VINP can make an ASR-effective DNN even more effective. Because VINP-oSpatialNet provides an anechoic speech prior with higher quality, its ASR performance is superior to VINP-TCN+SA+S. Notably, VINP-oSpatialNet achieves SOTA performance in ASR.
@@ IV-E2 Blind RIR Identification
Also setting the number of VEM iterations to 100, the RT60 and DRR estimation results are illustrated in Table III, where the bold font represents the best two results.

### v2

@@ IV-C2 Blind RIR Identification
Unprocessed | 1.48 | 0.70 | 2.37 | 3.20 | 13.1 | 5.6 | 4.6 | 1.31 | 2.82 | 24.1 | 7.9 | 5.7 |
Oracle | - | - | 3.76 | 3.90 | 7.6 | 4.5 | 4.0 | - | - | - | - | - |
GWPE [4] | 1.55 | 0.72 | 2.41 | 3.22 | 12.1 | 5.5 | 4.6 | 1.42 | 2.83 | 21.4 | 6.8 | 5.7 |
SkipConvNet [57] | 2.12 | 0.81 | 3.20 | 3.60 | 13.3 | 6.3 | 5.2 | 2.84 | 3.32 | 24.5 | 9.3 | 7.3 |
CMGAN [23] | 2.85 | 0.90 | 3.82 | 3.81 | 9.5 | 5.1 | 4.4 | 3.87 | 4.00 | 12.9 | 5.9 | 5.0 |
StoRM [25] | 2.34 | 0.86 | 3.73 | 3.96 | 11.4 | 6.2 | 5.1 | 3.72 | 4.01 | 17.5 | 10.2 | 8.0 |
@@
oSpatialNet* [53] | 2.87 | 0.92 | 3.57 | 3.88 | 8.9 | 4.9 | 4.3 | 3.48 | 3.87 | 10.5 | 5.4 | 4.5 |
VINP-TCN+SA+S (prop.) | 2.52 | 0.87 | 3.47 | 3.88 | 8.9 | 5.1 | 4.3 | 3.18 | 3.77 | 11.6 | 6.1 | 5.3 |
VINP-oSpatialNet (prop.) | 2.82 | 0.90 | 3.49 | 3.86 | 8.3 | 4.8 | 4.1 | 3.33 | 3.80 | 8.9 | 5.0 | 4.3 |
@@ IV-E1 Speech Dereverberation
Setting the number of VEM iterations to 100, the dereverberation results are presented in Table II, where the bold font denotes the best results. The results with negative dereverberation effects are marked with a darkened background.
As the metrics of unprocessed speech and oracle speech indicate, a larger ASR system implies a stronger recognition ability for oracle speech and stronger robustness to reverberation. Compared with using a large ASR system alone, introducing an ASR-effective front-end speech dereverberation system can significantly improve the performance of ASR with fewer parameters and lower computational costs. Among the comparison methods, only a few approaches are proven to be effective for ASR, such as GWPE, CMGAN, oSpatialNet*, and the proposed VINP-TCN+SA+S and VINP-oSpatialNet. The classical GWPE yields a marginal improvement in ASR performance. Among the DL-based methods, SkipConvNet, StoRM and TCN+SA+S encounter failures in ASR. TCN+SA+S and StoRM demonstrate marginal effectiveness solely when the ’tiny’ ASR model is employed. The reason for such ASR performance lies in the artifact induced by DNN, which is a common phenomenon in similar DL-based single-channel systems [26, 27]. Different from the comparison methods, VINP employs a linear CTF filtering process to model the reverberation effect and leverages the DNN output as a prior distribution of anechoic source speech. During the VBI procedure, VINP takes into account both this prior information and the observed data. By doing this, VINP circumvents the direct utilization of DNN output and thereby reduces the artifacts. Meanwhile, VINP still utilizes the powerful non-linear modeling ability of DNN. Consequently, VINP exhibits a remarkable superiority over the other approaches in ASR performance. When comparing VINP-TCN+SA+S with its backbone network TCN+SA+S, a gap in WER can be discerned, indicating that VINP can make an ASR-ineffective DNN effective. Both oSpatialNet* and VINP-oSpatialNet achieve ASR performance close to that of oracle speech on REVERB SimData when using the small and medium ASR systems. In other situations, a gap still exists, indicating that VINP can make an ASR-effective DNN even more effective. Because VINP-oSpatialNet provides an anechoic speech prior with higher quality, its ASR performance is superior to VINP-TCN+SA+S. Notably, VINP-oSpatialNet achieves SOTA performance in ASR.
@@ IV-E2 Blind RIR Identification
Also setting the number of VEM iterations to 100, the RT60 and DRR estimation results are illustrated in Table III, where the bold font represents the best results.